Calculate the GDP using EXPENDITURE APPROACH with the following hypothetical data:

Personal Consumption

3657

Depreciation

400

Wages

3254

Indirect Business Taxes

500

Interest

530

Domestic Investment

741

Government Expenditure

1098

Rental Income

17

Corporate Profits

744

Exports

673

Imports

704

Net Factor Payment

20









1
Expert's answer
2022-01-31T16:21:41-0500

Y = C + I + G + NX

Y = 3,657 + 741 + 1,098 + (673 – 704)

Y = 3,657 + 741 + 1,098 – 31

Y = 5,465